OpenStudy (anonymous):

The gravitational force, F, between an object and the Earth is inversely proportional to the square of the distance from the object to the center of the Earth. If an astronaut weighs 214 pounds on the surface of the Earth, what will this astronaut weigh 500 miles above the Earth? Assume that the radius of the Earth is 4000 miles.

OpenStudy (amistre64):

inversely suggests the constant of variation is a product: F(d^2) = k F = k/(d)^2

OpenStudy (amistre64):

k = 214/(4000)^2 k = ?/(4500)^2 equate the two and solve

OpenStudy (amistre64):

almost typed that correctly :/

OpenStudy (amistre64):

k = 214(4000)^2 k = ?(4500)^2 214(4000)^2 = ?(4500)^2 thats better

OpenStudy (anonymous):

I am still confused.

OpenStudy (amistre64):

your going to have to be more explicit than that since i cannot read minds ... yet :)

OpenStudy (anonymous):

would x be the question mark

OpenStudy (amistre64):

i spose you could use an "x", since it is conventionally considered to stand in the place of an unknown.

OpenStudy (amistre64):

i would have used an F since Force starts with an F